OpenStudy (anonymous):

The correct solution for the radical equation is: Select one: a. 5 and 10 b. –5 and –10 c. 5 d. 10

OpenStudy (anonymous):

\[\sqrt{x-1}=x+7\]

OpenStudy (anonymous):

okay so square both sides first

OpenStudy (anonymous):

guessm and check i guess 10\[\sqrt{10-1}=10+7?\\ \sqrt{9}=17?\\ 3=17\] no

OpenStudy (anonymous):

actually none of the choices are correct

OpenStudy (anonymous):

so what do i do

OpenStudy (anonymous):

check for a typo in your question

OpenStudy (anonymous):

there isnt one thats my actual question

OpenStudy (anonymous):

\[\sqrt{x-1}=x+7\] has no real number solutions

OpenStudy (anonymous):

then write them and tell them that there is a mistake in the question and you should get a free 100% on your next assignment for spotting this error

OpenStudy (anonymous):

they dont do that for me

OpenStudy (anonymous):

or at this school

OpenStudy (anonymous):

write them anyways there is no real solution to \[\sqrt{x-1}=x+7\] it is easy to check

OpenStudy (anonymous):

found it it was a

OpenStudy (anonymous):

@satellite73